Find all integer solutions m, n to the linear diophantine equation 174m + 54n = 18.
Answer #1

174 m+54n=18 ① g.c.d (174,54) = 6 174= 3(54 +12 54= 4812) + 6 12 = 2 x 6 +0 6=54- 4812 6 54-4 [174-3159 6 = 54 - 4(174) + 12(